A car starts from point A and reaches point B in 10 minutes after covering displacenient of 10 km. From point B car move to paint C in 15 minutes after covering displacement of 30 km. (a) Find average velocity
1
Expert's answer
2021-09-15T11:37:57-0400

"v_{avr}=(s_1+s_2)\/(t_1+t_2)=(10000+30000)\/(600+900)="


"="